OpenStudy (anonymous):

y=[x+(x+sin^2x)^5]^6

OpenStudy (anonymous):

i need to find the derivative

OpenStudy (anonymous):

y=[x+(x+sin^2x)^5]^6 y'=6(x+(x+sin^2x)^5)^5(1+5(x+sin^2x)^4(1+4sin2xcos2x))

OpenStudy (anonymous):

can you please explain to me what you did

OpenStudy (anonymous):

chain rule f(x)=u^6, f'(x)=6u^5u' u=[x+(x+sin^2x)^5] so f'(x)=6[x+(x+sin^2x)^5]^5[x+(x+sin^2x)^5]'

OpenStudy (anonymous):

[x+(x+sin^2x)^5]' need derivative again (1+5(x+sin^2x)^4(x+sin^2x)' this time (x+sin^2x)'=(1+2sinxcosx)=(1+sin2x)